# Reinforcement Learning Teachers of Test Time Scaling

### v2

@@ 1 Introduction
Figure 1: RLTs provide better student distillation and RL cold-starts than orders of magnitude larger LMs across competition and graduate-level tasks (AIME, MATH, GPQA).
Exploration is one of the critical challenges in reinforcement learning (RL) and has been a core focus of its literature [1, 2, 3]. Sparse rewards cannot yield any learning signal unless the agent is already capable of solving the given task at initialization. With the rise of RL for open-ended reasoning (RL reasoning) inducing a new form of language model (LM) scaling beyond prompt-engineering and search [4, 5], exploration has re-emerged as a key challenge. A canonical motivation for RL is the potential to bootstrap from partial solutions guided by the reward function and learn entirely new tasks from scratch. However, the nature of one-hot correctness rewards used in the RL reasoning framework fails to provide a dense form of guidance, focusing instead on reinforcing correct responses in the initial model’s pool of pass-at-k attempts – without true extrapolation beyond the LM’s initial latent abilities [6]. As a result, mostly large, already-capable models have been shown to improve consistently beyond cheaper and simpler supervised optimization [7].
@@
Based on these considerations, we propose a framework that avoids RL’s exploration challenge with a new class of specialized Reinforcement-Learned Teachers (RLTs) trained specifically to yield effective downstream distillation. Our main intuition is simple: the ability of real-world teachers is not measured by whether they can come up on their own with complex theorems, proofs, or answers from scratch. Instead, what matters is their ability to make use of readily available solutions and devise instructive explanations for their students. Thus, we depart from the traditional RL reasoning framework, tasking a model to first think and then come up with a new solution for the first time. Instead, RLTs are tasked with the easier problem of providing an effective explanation with the problem’s solution already given within their prompt. We train RLTs with dense rewards using the student’s log probabilities to assess its understanding of each problem’s ground-truth solution from the teacher’s explanations, and the interpretability of the logical leaps in the explanations themselves.
@@ 3.3 Evaluating the quality of explanations
    $r^{\mathrm{SS}}_{i}$: quantifying the student $\pi_{s}$ understanding of the solutions $s_{i}$ given the question $q_{i}$ and think tokens $t_{o_{i}}$ in context. This first reward term is computed with the student’s log probabilities over the solution tokens, reduced with both average and minimum operations:
@@
    $r^{\mathrm{KL}}_{i}$: quantifying whether the think tokens $t_{o_{i}}$ themselves are interpretable logical continuations from the student’s perspective as compared with the teacher’s. This second reward term is computed with the KL divergence over the same think tokens between the teacher’s distribution (under the RLT’s format with both $q_{i}$ and $s_{i}$ in context) and the student’s (with only the question $q_{i}$ in context), reduced with both average and maximum operations:
@@
Each term in our reward function serves a precise purpose. First, optimizing $r^{\mathrm{SS}}$ will produce explanations containing think tokens $t_{o_{i}}$ that maximize the student’s likelihood of reaching the correct solution $s_{i}$. However, this term alone does not differentiate between explanations that guide the student step-by-step and those that increase the solution’s likelihood without a logical path that can be learned from. An extreme instance of the latter would be an explanation simply repeating the solution tokens to increase likelihood, failing to provide general examples of reasoning methods that can be applied when approaching new problems. Thus, introducing $r^{\mathrm{KL}}$ fills precisely this gap, aligning the teacher’s distribution toward the student’s such that each think token in the output explanations cannot have too low probability when formatted in the distillation prompt $d_{i}$ with only the question $q_{i}$ and the previous think tokens in context. Intuitively, introducing this term regularizes for each step in the logical path traced by the teacher’s explanation to still make sense in the “student’s mind” given only its prior understanding and the question itself. Additionally, combining the average with min/max reductions ensures the rewards do not forego any individual token, regardless of the solution length or the number of think tokens in the teacher’s explanations. For instance, their omission could bias $r^{\mathrm{SS}}$ based on the length of the solutions or lead the teacher to prefer long explanations only to reduce the influence on $r^{\mathrm{KL}}$ of hard but necessary individual logical steps. For further discussion, we refer to Appendix 9, where we empirically analyze and validate all these design choices.
@@ 3.4 Pulling everything together: the RLT training paradigm
The RLT framework can be used with any RL algorithm (e.g., [17, 18]) with minimal modifications to the LM’s conditioning and reward, as described in the above subsections. In this work, we employ the simple GRPO recipe detailed in Section 2, resulting in the following training objective:
@@ 4.1 Training, distillation, and evaluation
We collect our distillation dataset with the learned RLTs using the same full set of 17K question-solution pairs from training. With the new reasoning traces, we then proceed to fine-tune our students either on this full data or a randomly sampled 1K subset, equating the distillation budget and following the same recipes as our baselines [8, 11]. Unlike previous RL distillation pipelines, we do not apply extra postprocessing refinements to improve the quality of the RLT’s reasoning traces, directly using our model’s raw outputs for student fine-tuning. We refer to Appendices A and B for further details regarding our training and distillation phases with complete lists of hyperparameters.
Following prior work [8, 12, 11], our main evaluation considers three popular and challenging tasks from the literature: AIME24 [23], the set of problems used for the American Invitational Mathematics Examination. MATH 500 [24], the set of problems selected by [25] from the canonical competition math benchmark. GPQA Diamond [26], the set of diamond difficulty problems on natural science topics from the Graduate-level Google-proof Q&A benchmark. We report the completion accuracy of each of our students using Lighteval [27]. When available, we use baseline results reported in prior work, which we found close to our early reproduction attempts. In Appendix C, we extend the experiments in this section by evaluating our models on additional tasks and settings.
@@ 4.2 Test-time reasoning across teachers and students
As shown in Table 1, the raw output explanations of our small 7B parameter RLT outperform all the considered data-distillation pipelines involving teachers with orders of magnitude more parameters and additional ad-hoc postprocessing steps. We also find the effectiveness of the RLT traces stays consistent across different data sizes, in contrast to the R1 traces from the Bespoke pipeline that appear significantly less effective when subsampled. Furthermore, even when distilling a Qwen-32B student, much larger than our 7B teacher, our RLT still outperforms all prior methods for both data sizes with considerable margins. We believe this result, in particular, shows how our framework could allow overcoming the current prohibitive costs of RL reasoning: shifting the burden of expensive RL procedures to small teachers, unable to effectively solve problems from scratch but highly specialized in the simpler task of producing effective explanations for large, more powerful students.
@@ 4.5 Explanation reward analysis
We also inspect qualitative examples chosen by selecting samples where the reward of the RLT explanations is particularly improved from the baseline R1 distillation pipeline. As shown in Figure 5, we find that the R1 traces with low rewards often try to rely on external tools, such as calculators, and employ language patterns likely idiosyncratic to the training data of the DeepSeek-V3 LM, as sentences with brief humorous comments [15]. Instead, the corresponding RLT explanations appear much more grounded and even manage to add new verification steps not considered by R1 to check the final solution. In Appendix 9, we provide additional examples showcasing further qualitative differences of our framework with R1 traces and also specific failure cases from training RLTs without proper balance between each reward component, such as repetitions and overly-long explanations.
@@ 5 Related work
A large part of the recent test-time scaling literature considering smaller LMs has focused on inducing reasoning with “teacher-student” supervised distillation [36], a widely validated technique in traditional LM development [22, 37]. This approach’s popularity to induce LM reasoning dates earlier than the RL paradigm, with older methods harnessing verifiers and prompting for self-improvement [38, 39]. By following a common structure of generation, filtering correct responses, and postprocessing them, modern RL-based distillation has seen significant advances mostly driven by more capable teachers [12, 11] and carefully curating targeted datasets [8, 16]. However, the effectiveness of current distillation pipelines was shown to be closely tied to the properties of the student itself [40, 9], and their ability to induce actual generalization remains unclear [41]. Unlike these traditional distillation pipelines, the RLT framework does not rely on verifiers for filtering, directly optimizes the teacher for downstream distillation, and does not require any post-processing, allowing direct transfer of reasoning capabilities to arbitrary tasks and even larger student models.
@@ 6 Discussion and extensions
This work introduced a new class of Reinforcement-Learned Teachers trained with a simpler dense-reward task that inputs both each problem’s question and solution, and optimizes the LM to provide instructive reasoning traces for distillation as outputs. Empirically, students trained or cold-started from the raw outputs of a 7B RLT obtain higher performance than prior distillation pipelines using orders-of-magnitude larger LMs. Furthermore, RLTs maintain their effectiveness even for distilling much larger students, and when providing reasoning traces for out-of-distribution tasks beyond their training corpus. Nonetheless, our work has only begun to study the design space of our new framework, with many exciting directions yet to be explored. One example is training RLTs and their students in tandem, allowing the teacher’s explanations to adapt to the student’s learning dynamics live. Pushing this further, the same model could even take both roles, iterating RL with our task formulation, providing access to each problem’s solution, to obtain instructive reasoning traces, and self-distillation [38] to revise its own explanation and learn how to solve questions from scratch – unifying the open-endedness of RL with the stability of supervised optimization.
Acknowledgments
The authors would like to thank Takuya Akiba for insightful project discussions, together with Yutaro Yamada and Boris Meinardus for providing feedback on early versions of the text.
@@ B.1 Reasoning traces generation
Learning rate $1\times 10^{-5}$ $1\times 10^{-5}$ $1\times 10^{-5}$
@@
Final learning rate $1\times 10^{-6}$ $1\times 10^{-6}$ $1\times 10^{-6}$
@@
Adam epsilon $1\times 10^{-8}$ $1\times 10^{-8}$ $1\times 10^{-8}$
@@ B.2 Student distillation specifications
Learning rate $1\times 10^{-6}$ $1\times 10^{-6}$
@@
Final learning rate $1\times 10^{-6}$ $1\times 10^{-6}$
@@
Adam epsilon $1\times 10^{-8}$ $1\times 10^{-8}$

### v3

@@ 1 Introduction
Figure 1: RLTs provide better student distillation and RL cold-starts than orders of magnitude larger LMs across competition and graduate-level tasks (AIME, MATH, GPQA). This holds when distilling students of the same size (Left) and also 32B students, much larger than the RLT itself (Right).
Exploration is one of the critical challenges in reinforcement learning (RL) and has been a core focus of its literature [1, 2, 3]. Sparse rewards cannot yield any learning signal unless the agent is already capable of solving the given task at initialization. With the rise of RL for open-ended reasoning (RL reasoning) inducing a new form of language model (LM) scaling [4, 5, 6] beyond prompt-engineering and search [7, 8], exploration has re-emerged as a key challenge. A canonical motivation for RL is the potential to bootstrap from partial solutions, guided by an informative reward function, and learn entirely new tasks from scratch. However, the nature of one-hot correctness rewards used in the RL reasoning framework fails to provide a dense form of guidance to assess relative progress, focusing instead on reinforcing correct responses in the initial model’s pool of pass-at-k attempts – without true extrapolation beyond the LM’s initial latent abilities [9]. As a result, mostly large, already-capable models have been shown to improve consistently beyond cheaper and simpler supervised optimization [4].
@@
Based on these considerations, we propose a framework that avoids RL’s exploration challenge with a new class of specialized Reinforcement-Learned Teachers (RLTs) trained specifically to yield effective downstream distillation. Our main intuition is simple: the ability of real-world teachers is not measured by whether they can come up on their own with complex theorems, proofs, or answers from scratch. Instead, what matters is their ability to make use of readily available solutions and devise instructive explanations for their students. Thus, we depart from the traditional RL reasoning framework, tasking a model to first think and then come up with a new solution for the first time. Instead, RLTs are tasked with the easier problem of providing an effective explanation with the problem’s solution already given within their prompt. We train RLTs with dense and informative rewards obtained from the student’s log probabilities. These rewards provide an intuitive measure of the student’s understanding of each problem’s ground-truth solution following the teacher’s explanations, and the interpretability of the logical leaps in the explanations themselves.
@@ 3.3 Evaluating the quality of explanations
    $r^{\mathrm{SS}}_{i}$: quantifying the student $\pi_{s}$ understanding of the solutions $s_{i}$ given the question $q_{i}$ and think tokens $t_{o_{i}}$ in context. This first reward term is computed with the student’s log probabilities over the solution tokens. We reduce this vector to a scalar by applying both average and minimum operations over the different per-token values:
@@
    $r^{\mathrm{KL}}_{i}$: quantifying whether the think tokens $t_{o_{i}}$ themselves are interpretable logical continuations from the student’s perspective as compared with the teacher’s. This second reward term is computed with the KL divergence over the same think tokens between the teacher’s distribution (under the RLT’s format with both $q_{i}$ and $s_{i}$ in context) and the student’s (with only the question $q_{i}$ in context). Similarly to $r^{\mathrm{SS}}_{i}$, we reduce this vector to a scalar by applying both average and maximum operations over the different per-token values:
@@
Each term in our reward function serves a precise purpose. First, optimizing $r^{\mathrm{SS}}$ will produce explanations containing think tokens $t_{o_{i}}$ that maximize the student’s likelihood of reaching the correct solution $s_{i}$. However, this term alone does not differentiate between explanations that guide the student step-by-step and those that increase the solution’s likelihood without a logical path that can be learned from. An extreme instance of the latter would be an explanation simply repeating the solution tokens to increase likelihood, failing to provide the student with general examples of reasoning methods that can be applied when approaching new problems. Thus, introducing $r^{\mathrm{KL}}$ fills precisely this gap, aligning the teacher’s distribution toward the student’s such that each think token in the output explanations cannot have too low probability when formatted in the distillation prompt $d_{i}$ with only the question $q_{i}$ and the previous think tokens in context.
Intuitively, introducing this term regularizes for each step in the logical path traced by the teacher’s explanation to still make sense in the “student’s mind” given only its prior understanding and the question itself. We note that if we instead compared two distributions conditioned on the question alone, the KL would vanish, and $r^{KL}$ would fail to penalize the RLT steps that the teacher could not have generated without the solution in context. Additionally, combining the average with min/max reduction terms ensures the rewards do not forego any individual token, regardless of the solution length or the number of think tokens in the teacher’s explanations. For instance, their omission could bias $r^{\mathrm{SS}}$ based on the length of the solutions or lead the teacher to prefer long explanations only to reduce the influence on $r^{\mathrm{KL}}$ of hard but necessary individual logical steps. For further discussion, we refer to Appendix 14, where we empirically analyze and validate each design choice in our reward function in terms of final performance (Table 14) and concrete qualitative differences of the resulting explanations (Figures 8 through 12).
@@ 3.4 Pulling everything together: the RLT training paradigm
The RLT framework can be used with any RL algorithm (e.g., [18, 19]) with minimal modifications to the LM’s conditioning and reward, as described in the above subsections. For all our main experiments, we employ the simple GRPO recipe detailed in Section 2, resulting in the following training objective:
@@ 4.1 Training, distillation, and evaluation
We collect our distillation dataset with the learned RLTs using the same full set of 17K question-solution pairs from training. With the new reasoning traces, we then proceed to fine-tune our students either on this full data or a randomly sampled 1K subset, equating the distillation budget and following the same recipes as our baselines [6, 12]. Unlike previous RL distillation pipelines, we do not apply extra postprocessing refinements to improve the quality of the RLT’s reasoning traces, directly using our model’s raw outputs for student fine-tuning. We note this is in contrast to prior distillation pipelines that make use of the ground-truth answers to verify the correctness of their reasoning traces and rely on multi-generations, filtering, and post-processing stages to improve data quality [24, 13]. We refer to Appendices A and B for further details regarding our training and distillation phases with complete lists of hyperparameters.
Following prior work [6, 13, 12], our main evaluation considers three popular and challenging tasks from the literature: AIME24 [25], the set of problems used for the American Invitational Mathematics Examination. MATH 500 [26], the set of problems selected by [27] from the canonical competition math benchmark. GPQA Diamond [28], the set of diamond difficulty problems on natural science topics from the Graduate-level Google-proof Q&A benchmark. We report the completion accuracy of each of our students using Lighteval [29]. When available, we use baseline results reported in prior work, which we found close to our early reproduction attempts. In Appendix C, we extend the experiments in this section by evaluating our models on additional tasks and analyzing the impact of teacher size, the base RL algorithm, and the student used during training.
@@ 4.2 Test-time reasoning across teachers and students
As shown in Table 1, the raw output explanations of our small 7B parameter RLT outperform all the considered data-distillation pipelines involving teachers with orders of magnitude more parameters and additional ad-hoc postprocessing steps. We also find that the effectiveness of the RLT traces stays consistent across different data sizes. In contrast, the R1 traces from the Bespoke pipeline appear significantly less effective when subsampled. Furthermore, even when distilling a Qwen-32B student, much larger than our 7B teacher, our RLT still outperforms all prior methods for both data sizes with considerable margins. We believe this result, in particular, shows how our framework could allow overcoming the current prohibitive costs of RL reasoning: shifting the burden of expensive RL procedures to small teachers, unable to effectively solve problems from scratch but highly specialized in the simpler task of producing effective explanations for large, more powerful students.
@@ 4.5 Explanation reward analysis
We also inspect qualitative examples chosen by selecting samples where the reward of the RLT explanations is particularly improved from the baseline R1 distillation pipeline. As shown in Figure 5, we find that the R1 traces with low rewards often try to rely on external tools, such as calculators, and employ language patterns likely idiosyncratic to the training data of the DeepSeek-V3 LM, as sentences with brief humorous comments [16]. Instead, the corresponding RLT explanations appear much more grounded and even manage to add new alternative verification steps not considered by R1 to check the final solution. In Appendix 14, we provide additional examples showcasing further qualitative differences of our framework with R1 traces and also specific failure cases from training RLTs without proper balance between each reward component, such as repetitions and overly-long explanations.
@@ 5 Related work
A large part of the recent test-time scaling literature considering smaller LMs has focused on inducing reasoning with “teacher-student” supervised distillation [36], a widely validated technique in traditional LM development [23, 37]. This approach’s popularity to induce LM reasoning dates earlier than the RL paradigm, with older methods harnessing verifiers and prompting for self-improvement [38, 39]. Part of this earlier exploration showed preliminary signs of how teachers could be considerably steered to provide better student data [40, 41] and how improvements in teaching also led to improvements in question-answering domains [42]. Lately, by following a common structure of generation, filtering correct responses, and postprocessing them, modern RL-based distillation has seen significant advances mostly driven by more capable teachers [13, 12] and carefully curating targeted datasets [6, 17]. However, the effectiveness of current distillation pipelines was shown to be closely tied to the properties of the student itself [43, 10], and their ability to induce actual generalization remains unclear [44]. Unlike these traditional distillation pipelines, the RLT framework does not rely on verifiers for filtering, directly optimizes the teacher for downstream distillation, and does not require any post-processing, allowing direct transfer of reasoning capabilities to arbitrary tasks and even larger student models.
@@ 6 Discussion and extensions
This work introduced a new class of Reinforcement-Learned Teachers trained with a simpler dense-reward task that inputs both each problem’s question and solution, and optimizes the LM to provide instructive reasoning traces for distillation as outputs. Empirically, students trained or cold-started from the raw outputs of a 7B RLT obtain higher performance than prior distillation pipelines using orders-of-magnitude larger LMs. Furthermore, RLTs maintain their effectiveness also when providing reasoning traces for out-of-distribution tasks beyond their training corpus, and even for distilling much larger students than the teacher itself. These remarkable properties can have significant implications for the scalability of developing reasoning models, due to the disproportionate cost of RL compared to all other post-training stages: while distilling a 32B student on fixed traces took less than a day on a single H100 compute node, training this larger model using RL on the same questions would have taken months with the same hardware. Nonetheless, our work has only begun to study the design space of our new framework, with many exciting directions yet to be explored. One example is training RLTs and their students in tandem, allowing the teacher’s explanations to adapt to the student’s learning dynamics live. Pushing this further, the same model could even take both roles, iterating RL with our task formulation, providing access to each problem’s solution, to obtain instructive step-by-step reasoning traces, and self-distillation [38] to revise its own explanation and learn how to solve questions from scratch – unifying the open-endedness of RL with the stability of supervised optimization.
@@ A.2 Traditional reasoning and RLT formats
A.3 Cost comparison
Table 4: Cost to train and generate datasets with RLT models as opposed to DeepSeek R1 [4].
Model/GPU-hours Training (GPU model) Data generation (GPU model)
DeepSeek R1 >688000 (H800) >1067 (H100)
7B RLT teacher 280.4 (H100) 6.7 (H100)
In Table 4, we report the GPU cost of training and collecting distillation data with our RLT. We compared these times with an estimated cost of traditional RL post-training and data collection using the DeepSeek R1 model. To compute this estimate, we used official figures for the GPUs used to train this DeepSeek model (2,048 NVIDIA H800 GPUs [16]) and online communication from the DeepSeek team for the training time. We calculated a lower bound of the GPU hours for generating the distillation dataset by multiplying the minimum amount of GPU resources required to host DeepSeek R1 with half precision and the “roof inference speed” for generating each completion in our dataset, without taking into account slowdowns from the growing context size. However, we note that both these estimates should be considered large approximations as we did not even consider time/resources for multiple generations, filtering, and post-processing stages conducted in our baselines [12, 13].
@@ B.1 Reasoning traces generation
Learning rate $\times 10^{-5} 1×1$$\times 10^{-5}$ $1\times 10^{-5}$
@@
Final learning rate $\times 10^{-6} 1×1$$\times 10^{-6}$ $1\times 10^{-6}$
@@
Adam epsilon $\times 10^{-8} 1×1$$\times 10^{-8}$ $1\times 10^{-8}$
@@ B.2 Student distillation specifications
Learning rate $\times 10^{-6} 1×1$$\times 10^{-6}$
@@
Final learning rate $\times 10^{-6} 1×1$$\times 10^{-6}$
@@
Adam epsilon $\times 10^{-8} 1×1$$\times 10^{-8}$
@@ C.1 Coding and multilingual reasoning
C.2 The generality of RLT across RL algorithms
Table 10: Student performance from RLTs trained with GRPO [15], and RLOO [19].
Evaluated model Data size AIME 2024 MATH 500 GPQA Diamond Overall
Qwen2.5-7B-Instruct N/A 10.00 74.20 33.30 39.20
Bespoke-7B 17K 20.00 82.00 37.80 46.60
RLT-7B 17K 23.30 82.80 42.40 49.50
RLT-7B (RLOO) 17K 20.00 83.60 42.90 48.80
To validate the generality of the RLT framework, we extended our implementation by training our 7B RLT model using the RLOO [19] algorithm instead of GRPO. We provide this comparison in Table 10 above. In line with recent empirical findings about the similar empirical effectiveness of the different reasoning RL algorithms [51], the performance of our RLT trained with RLOO appears very close to our results obtained using GRPO in the main text. While there exists a small gap, we believe this is mostly due to the fact that we did not re-adjust any of the main hyperparameters from our GRPO implementation (e.g., batch size, number of generations per question). These results provide further concrete evidence confirming that the RLT paradigm’s performance is not tied to any specific RL algorithm.
C.3 The effects of teacher scale
Table 11: Student performance from RLTs with 3B and 7B parameters.
Evaluated model Data size AIME 2024 MATH 500 GPQA Diamond Overall
Qwen2.5-7B-Instruct N/A 10.00 74.20 33.30 39.20
RLT-7B 17K 23.30 82.80 42.40 49.50
RLT-7B (3B teacher) 17K 20.00 80.60 38.90 46.50
Qwen2.5-32B-Instruct N/A 26.70 84.00 49.00 53.20
RLT-32B 17K 66.70 93.40 59.60 73.20
RLT-32B (3B teacher) 17K 46.70 91.40 53.50 63.90
To validate the effects of the RLT teacher’s scale, we extended our implementation by training an RLT model with 3B parameters. The only change to our hyperparameters was to double the number of supervised finetuning iterations, but kept the same number of RL training steps. As shown in Table 11, we expectedly find that larger teachers yield better explanations and downstream results. However, we find that most of the performance gap between our 3B and 7B teachers occurs when distilling the 32B student LM, where the mismatch between teacher and student capabilities tends to the extreme. However, even in this case where the student is over 10x larger than the teacher, our 3B RLT is still able to provide considerable improvements to the initial Qwen 32B student. The reason that we did not observe performance degradation appears to be due to the role of the $r^{KL}$ reward term plays during the optimization. Even in cases where the initial 3B teacher is unable to provide logical explanations to a specific question and answer, optimizing $r^{KL}$ will naturally guide the teacher’s output distribution (with both question and answer in context) toward converging to what was most likely from the student’s distribution (with only the question in context).
C.4 Stronger students make stronger teachers
Table 12: Student performance using RLTs trained with a combination of 7B and 32B students, and following a 7B student updated midway through training.
Evaluated model Data size AIME 2024 MATH 500 GPQA Diamond Overall
Qwen2.5-7B-Instruct N/A 10.00 74.20 33.30 39.20
Bespoke-7B 17K 20.00 82.00 37.80 46.60
RLT-7B 17K 23.30 82.80 42.40 49.50
RLT-7B (7B + 32B) 17K 23.30 83.80 41.90 49.70
RLT-7B (2-stage) 17K 26.70 84.00 41.40 50.70
We extended our implementation to analyze how scaling up student capabilities to compute the terms in the RLT reward affects the training of our new models. First, we considered ensembling both the 7B and the 32B students, averaging their probabilities, thus also introducing a degree of diversity into the RLT reward. Second, we considered conducting training in two stages, each optimizing both teacher and student for half the total number of optimization steps. In particular, we paused the teacher RL optimization midway, distilled a student with this preliminary RLT’s explanations, and resumed our teacher optimization for the remaining steps with this updated student to compute the RLT reward. This second extension has the added benefit of better aligning the teacher rewards to the student’s learning dynamics, and can be considered an initial step toward online optimization of both models, which we believe to be an interesting future direction, as outlined in Section 6. As shown in Table 12 above, both these extensions provide noticeable improvements to the RLT distillation performance. Overall, these results, together with the ones showing the effects of teacher scale from Table 11, further highlight the complementarity of our new framework to harness future models and LLM advances across different beneficial axes.
C.5 Student performance and scale of RLT dataset
Table 13: The effects of data size from training with RLT-generated data from different numbers of questions from Li et al. [12] (1K and 17K) and its NuminaMath split [52].
Evaluated model Data size AIME 2024 MATH 500 GPQA Diamond Overall
Qwen2.5-7B-Instruct N/A 10.00 74.20 33.30 39.20
Bespoke-7B-1K 1K 13.30 80.00 33.80 42.40
Bespoke-7B 17K 20.00 82.00 37.80 46.60
Bespoke-7B (1K NuminaMath) 1K 16.70 79.00 33.30 43.00
Bespoke-7B (10K NuminaMath) 10K 16.70 79.60 40.90 45.70
RLT-7B-1K 1K 20.00 80.40 41.90 47.40
RLT-7B 17K 23.30 82.80 42.40 49.50
RLT-7B (1K NuminaMath) 1K 16.70 80.40 40.40 45.80
RLT-7B (10K NuminaMath) 10K 20.00 81.00 43.40 48.10
We collected additional results by changing the amount of student distillation data, only focusing on the NuminaMath dataset [52]. We constructed new datasets with 1K and 10K samples using the AIME, MATH, and Olympiads subsets of NuminaMath from Li et al. [12], enabling us to directly compare with the available reasoning traces from our best baseline using postprocessed reasoning traces from DeepSeek R1, used for Bespoke-7B. As shown in the Table 13 above, we find that the performance of the RLT and R1 traces from the Bespoke pipeline positively scales with the amount of student training data. While we removed all science questions from STILL-2 present in our original data, we find performance on GPQA does not decrease substantially, evidencing positive transfer across different reasoning domains. We note these results are consistent with our previous random subsampling of the questions from Li et al. [12], with the RLT traces outperforming the postprocessed R1 traces across all tested data sizes and configurations.